OpenStudy (anonymous):

i am not sure f this is correct can someone please check Solve 6x = 24 x ≈ 30.84 x ≈ 1.77 x ≈ 0.56 x ≈ 0.23--->My Answer

OpenStudy (anonymous):

Just Divide 24 By 6. And You Got It.

OpenStudy (anonymous):

but that isnt one of my choices

OpenStudy (anonymous):

because that is what i did first

OpenStudy (anonymous):

divide 24 by 6

OpenStudy (anonymous):

I think you meant 6^x = 24

OpenStudy (anonymous):

yes i thought i put the carrot

OpenStudy (anonymous):

but i didnt sorry about that

OpenStudy (anonymous):

Ohhh Okay.

OpenStudy (anonymous):

a^b = c b = log_a c = lg c / lg a = ln c / ln a

OpenStudy (anonymous):

u lost me i am not good with this stuff

OpenStudy (anonymous):

a^x = c x = lg c / lg a 6^x = 24 In this case, a = 6 and c = 24.